OpenStudy (anonymous):

solve sqrt 5x-8=4 please 10 minutes left!

OpenStudy (anonymous):

\[\sqrt{(5x - 8)} = 4\] Is it the question???

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Try squaring both the sides, what will you get can you answer me???

OpenStudy (anonymous):

is it 25x-56?

OpenStudy (anonymous):

No rachel.. what is the square of :\[\sqrt{(5x - 8)}\] ?? Tell me after squaring this what will you get??

OpenStudy (anonymous):

so im squaring the 5 and the 8?

OpenStudy (anonymous):

No dear, it is already in square root.. Remember, square of : \[\sqrt{x} = x\] i.e: \[(\sqrt{x})^2 = x\] Getting or not??

OpenStudy (anonymous):

no

OpenStudy (anonymous):

what is the point you are not getting??

OpenStudy (anonymous):

the squaring part

OpenStudy (anonymous):

See, tell me what is the square of 5??

OpenStudy (anonymous):

25

OpenStudy (anonymous):

May I write is as : \[5^2 = 5 \times 5\] Yes or not??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

If instead of 5, it would be \[\sqrt{5}\] Then how will I write???

OpenStudy (anonymous):

5?

OpenStudy (anonymous):

\[\sqrt{5}^2 = \sqrt{5} \times \sqrt{5}\]

OpenStudy (anonymous):

\[= \sqrt{5}\]

OpenStudy (anonymous):

whats next??

OpenStudy (anonymous):

So,what is the square of: \[\sqrt{(5x - 8)}\]???